Indira paid interest charges of $189.93 on a $4,870 invoice that was two months overdue.

What monthly rate of simple interest was she charged?

Answer by TimothyLamb(4379)

compound-interest formula:
v = p( 1 + (r/n) )^(nt)
where:
v = final value
p = initial (principal) value
r = interest rate (as a decimal)
n = number of compounding periods per year
t = number of years
---
get the interest amount:
i = v - p
i = p( 1 + (r/n) )^(nt) - p
189.93 = 4870*( 1 + (r/1) )^(1*2) - 4870
189.93 = 4870*( 1 + r )^2 - 4870
4870*( 1 + r )^2 = 4870 + 189.93
( 1 + r )^2 = (4870 + 189.93)/4870
sqrt( ( 1 + r )^2 ) = sqrt( (4870 + 189.93)/4870 )
1 + r = sqrt( (4870 + 189.93)/4870 )
r = sqrt( (4870 + 189.93)/4870 ) - 1
r = 0.0193134944658
---
monthly rate = 0.0193134944658 * 100
monthly rate = 1.93135%
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---